Ask your own question, for FREE!
Mathematics 9 Online
OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 3 or a 5 on the second die, given that you rolled an odd number on the first die? A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

@oldrin.bataku @Euler271

OpenStudy (anonymous):

the outcomes are independent of each other therefore, the first die doesn't matter. eveyr face still has 1/6 chance. so 3 OR 5 is 1/6 + 1/6 = 1/3

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

thnks!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
20 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
23 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!